Leblanc and the Development of Twentieth Century Logic

Lisa O’Neill & John Serembus

Twentieth century logic witnessed the development of alternative logics, among them logics that sought to address philosophical problems more concretely than classical approaches. Within these alternatives are two interesting and important areas: free logic and probability logic. A key figure in the development of these fields was logician and philosopher Hugues Leblanc (1924-1999).

Along with others Leblanc developed free logic, an alternative to classical logic and a kind of logic that claims to address certain philosophical problems more directly and extensively than classical logic. Additionally, he was a pioneer in the development of probability logic, in particular, probabilistic semantics. Probabilistic semantics is an alternative to standard model-theoretic semantics and an approach to semantics that also claims to boast useful philosophical solutions. In recognition of his numerous contributions we present a paper that focuses on Leblanc’s most important contributions in these two areas. Specific attention will be given to the usefulness of these alternatives in addressing key philosophical problems.

Addresses Lisa O’Neill: Department of Philosophy, Saint Peter’s College, 2641 Kennedy Blvd., Jersey City, NJ 07306, USA

E-mail: LAONEILL@EROLS.COM

John Serembus: Humanities Division, Widener University, Chester, PA 19013, USA